OpenStudy (anonymous):

Part 1. Using the two functions listed below, insert numbers in place of the letters a, b, c, and d so that f(x) and g(x) are inverses. f(x)= x+a b g(x)=cx−d Part 2. Show your work to prove that the inverse of f(x) is g(x). Part 3. Show your work to evaluate g(f(x)). Part 4. Graph your two functions on a coordinate plane. Include a table of values for each function. Include 5 values for each function. Graph the line y = x on the same graph.

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@SedateFrog712

OpenStudy (anonymous):

@aum

OpenStudy (aum):

f(x) = x + a ??? 'b' is all by itself on a separate line. Where does it belong?

OpenStudy (anonymous):

sorry thought i fixed it already \[f(x)=\frac{ x+a }{ b }\]

OpenStudy (aum):

Find the inverse of \[f(x)=\frac{ x+a }{ b } \\ y = \frac{ x+a }{ b } \\ \text{Switch x and y:} \\ x = \frac{ y+a }{ b } \\ \text{Solve for y:} \\ x * b = y + a \\ y = bx - a \\ f^{-1}(x) = bx - a \\ \text{Compare it to g(x): } \\ g(x) = cx - d \\ b = c \\ a = d \]

OpenStudy (anonymous):

followed you til we got to g(x)

OpenStudy (aum):

The problem states g(x) is the inverse function and that g(x) = cx - d

OpenStudy (aum):

In order for g(x) to be the inverse of \(f^{-1}(x)\), b must equal c and a must equal d. Pick any values for c and d. Example c = 2; d = 3. Then a = 3 and b = 2 Therefore, \(\large f(x)= \Large \frac{ x+3 }{ 2 } \) and \(\large g(x) = 2x - 3\) are inverses.

OpenStudy (aum):

Part 2) To prove g(x) is the inverse of f(x), find f(g(x)) and prove it is x. If f(g(x)) = x then g(x) is the inverse of f(x).

OpenStudy (aum):

Part 3) Evaluate g(f(x)) Part 4) Both f(x) and g(x) are straight lines. Just two points on each line will be sufficient to graph each line. But they want a table of five x values for each function. You can choose x = -2, -1, 0, 1, 2 and calculate f(x) and g(x) and put them in a table. The line y = x will be a 45 degree straight line passing through the origin. Since f(x) and g(x) are inverse of each other, their graphs should be symmetric with respect to the y = x line.